OpenStudy (anonymous):

I have to find delta given that as x approaches 5, 1/x=1/5. The epsilon is 0.05. The correct answer 1/505, but I'm not sure how to get that answer.

OpenStudy (aum):

Double check your answer. The correct answer is delta = 1.  f(x) = 1/x f(a) = 1/a = 0.25 = 1/4. Therefore, a = 4. f(b) = 1/b = 0.15 = 3/20. Therefore, b = 20/3 Find the minimum of |b - 5| and |5 - a| : Minimum of |20/3 - 5| = 5/3 and |5 - 4| = 1 is 1. Minimum is 1. So delta = 1.